OpenStudy (anonymous):

find the equation of the tangent for the graph y=ln x at the point where y=1

OpenStudy (demandandsupply):

Differentiate y=lnx

OpenStudy (anonymous):

i did that and got a slope of 1/x and then compared that to 1 to find what the slope is when y=1, and got 1. i then compared 1 to (ln x) to find what x is when y =1 and got e^1. i now have a coordinate and a slope and plugging that into m(x-x1)=y-y1 i got y = 2 - e/x which is not the correct answer. where did i go wrong? the correct answer is y = x/e

OpenStudy (demandandsupply):

You are almost there, most of that is right :) Give me one moment

OpenStudy (demandandsupply):

You're x coordinate should be e, y=1, and equation for gradient is, 1/x

OpenStudy (demandandsupply):

substitute your x value in 1/x, to get 1/e

OpenStudy (demandandsupply):

y-y1=m(x-x1). Substitute your values in

OpenStudy (demandandsupply):

y-1=1/e(x-e)

OpenStudy (demandandsupply):

Times everything by e - ey-e=x-e

OpenStudy (demandandsupply):

Rearrange for the subject with 'y' in it. ey=x

OpenStudy (demandandsupply):

Then, divide both sides by e and bam. y=e/x :)

OpenStudy (anonymous):

thanks so much! I got it!

OpenStudy (anonymous):

just a note: instead of multiplying everything by e i expanded which brings to the same results.

OpenStudy (demandandsupply):

No problem :) Yes that works too, the first way is just how ive always done it because you dont have to deal with fractions which can be a bit tricky

OpenStudy (anonymous):

Right. Thanks again!